# A Novel Compact LLM Framework for Local, High-Privacy EHR Data Applications

### v1

@@ 3.4 LLM fine tuning
To enhance the model’s adaptability and responsiveness to our specific task without extensively altering its pre-trained weights, we incorporated the Low-Rank Adaptation (LoRA, [32]) technique. LoRA allows us to dynamically adjust the model’s behavior by introducing trainable parameters that focus on specific parts of the model, such as attention and feed-forward layers. The trainable parameters is around 3.04% for Gemma-2b and 2.29% for Gemma-7b.
@@
Once the model has been thoroughly trained and validated, we apply it to the entire dataset, not limiting its use to only the samples involved in the training process. We then utilize the outcomes from this comprehensive application for subsequent validation.
Due to privacy restrictions, we are not permitted to use the Private HNC Dataset for model fine-tuning. As a result, we exclusively use the MIMIC-IV dataset for fine-tuning the model.
@@ 4.1 Introduction of Two EHR Datasets
The MIMIC-IV is an extensive, anonymized EHR database encompassing data from over 70,000 patients admitted to Beth Israel Deaconess Medical Center between 2008 and 2019 [9]. This period covers more than a decade of intensive care unit (ICU) admissions, making it a pivotal resource for medical research. With 26 tables of detailed patient data including demographics, lab measurements, procedures, medications, ICD codes, and clinical notes, MIMIC-IV offers comprehensive access to real-world clinical data. This database adheres strictly to the Health Insurance Portability and Accountability Act (HIPAA), ensuring patient privacy while supporting a wide range of medical studies. The breadth and depth of MIMIC-IV facilitate large-scale analyses critical for advancing clinical practice and shaping health policy.
We filter for subjects with at least one clinical note containing metastasis-related keywords, resulting in a dataset comprising notes from 13,923 unique subjects across 28,390 hospital admissions. After examining the ICD codes, 8,510 (approximately 61.1%) of these subjects have ICD-9 codes 197, 198, 199, or ICD-10 codes C78, C79, C80, representing metastasis of tumor. Additionally, 18,552 hospital admissions (approximately 65.3%) are associated with these ICD codes.
@@ 4.2 Evaluation metrics
“Time-sensitive” accuracy defined on the Private HNC dataset
Creating a definitive gold standard for evaluating experimental outcomes in this study presents a complex challenge, primarily due to the incomplete reflection of diseases through ICD codes alone. To address this and ensure a robust comparative analysis, we employ a methodological approach for our Private HNC Dataset centered around the use of ICD-10 codes C78, C79, and C80, which are indicative of metastatic conditions. It is crucial to acknowledge, however, that the absence of these codes does not unequivocally rule out the presence of metastasis in a patient.
For each patient diagnosed with metastasis according to the aforementioned codes, we review their clinical notes over a 20-day period, encompassing 10 days before and after the diagnosis date. This time frame is selected to capture the most relevant clinical observations related to the diagnosis. During this period, we conduct a systematic search for predefined keywords associated with metastasis in the clinical notes, utilizing both zero-shot and few-shot learning, with or without the preprocessing step.
The classification accuracy of our model is assessed based on the relative counts of notes: if the number of notes indicating metastasis exceeds those indicating its absence, the classification is considered correct. Conversely, if non-metastatic notes predominate, the classification is considered incorrect. Cases with an equal count are classified as inconclusive due to insufficient information. We denote the total number of correct, incorrect, and inconclusive classifications as $N_{c}$, $N_{i}$, and $N_{l}$, respectively. Additionally, if a review period lacks clinical notes altogether, we interpret this as an absence of relevant information rather than a misclassification by the LLM, and these instances are consequently excluded from our analysis. We use $P_{c}=\frac{N_{c}}{N_{c}+N_{i}+N_{l}}$ to denote the proportion of correct classifications out of all evaluated cases. Similarly, we use $P_{i}=\frac{N_{i}}{N_{c}+N_{i}+N_{l}}$ and $P_{l}=\frac{N_{l}}{N_{c}+N_{i}+N_{l}}$ to denote the proportions of incorrect and inconclusive classifications, respectively.
@@
For the MIMIC-IV dataset, the absence of timestamps for specific ICD codes precludes us from employing the same evaluative approach as with the Private HNC Dataset. Consequently, we adapt our methodology to assess the incidence of metastasis among patients through alternative means. Our modified process begins with the exclusion of all clinical notes lacking keywords pertinent to metastasis. Subsequently, we deploy both zero-shot and few-shot LLMs, applying these with and without our preprocessing techniques, to analyze the remaining notes. We also use fine-tuned LLM to classify the clinical notes.
@@
For each patient associated with ICD-9 codes 197, 198, 199, or ICD-10 codes C78, C79, or C80, we conduct a detailed analysis by tallying instances where the LLMs have classified clinical notes as indicative of metastasis. A patient is considered correctly classified if the number of notes flagged as metastatic exceeds those identified as non-metastatic. This method provides a focused evaluation of metastatic conditions within the dataset. To simplify the notation, we let $M_{LLM}^{i}$ and $N_{LLM}^{i}$ denote the group of patients classified as having metastasis and non-metastasis, respectively, by the LLM. Similarly, we let $M_{ICD}^{i}$ and $N_{ICD}^{i}$ represent the group of patients with and without metastasis based on the ICD codes, respectively. We then calculate #$\frac{\#(M_{LLM}^{i}\cap M_{ICD}^{i})}{\#(M_{ICD}^{i})}$ to denote the proportion of patients correctly identified as having metastasis by the LLM out of all patients with a metastatic diagnosis according to ICD codes, where #()#\#($\cdot$)# ( ⋅ ) denotes the cardinality (number of elements) of a set.
@@
We extend this approach to individual hospital admissions, allowing for a more granular assessment of the LLMs’ performance across different patient encounters. We use $M_{LLM}^{ha}$ and $N_{LLM}^{ha}$ to denote the group of hospital admissions classified as having metastasis and non-metastasis, respectively, by the LLM. Similarly, $M_{ICD}^{ha}$ and $N_{ICD}^{ha}$ represent the group of admissions with and without metastasis based on the ICD codes, respectively. We then calculate #$\frac{\#(M_{LLM}^{ha}\cap M_{ICD}^{ha})}{\#(M_{ICD}^{ha})}$ to denote the proportion of hospital admissions correctly identified as having metastasis by the LLM out of all admissions diagnosed with metastasis according to ICD codes.
@@
The analysis above serves as a measure of “so-called” sensitivity. It is crucial to recognize that even an ideal model is unlikely to achieve 100% sensitivity, as clinicians may not always explicitly document metastasis in clinical notes, despite assigning a corresponding ICD code. Nevertheless, the observed trends offer valuable insights into the model’s performance and its effectiveness in these complex scenarios. Conversely, the absence of an ICD code does not necessarily indicate that metastasis is not mentioned in the notes. Ideally, for patients lacking the relevant ICD codes, the LLMs should more frequently classify them as either not having metastasis or lacking sufficient information. To assess this, we calculate the “so-called” specificity as the proportion of patients without the relevant ICD codes who are correctly classified by the LLMs as non-metastatic or inconclusive, represented by #$\frac{\#(N_{LLM}^{i}\cap N_{ICD}^{i})}{\#(N_{ICD}^{i})}$ for individual patients and #$\frac{\#(N_{LLM}^{ha}\cap N_{ICD}^{ha})}{\#(N_{ICD}^{ha})}$ for hospital admissions.

### v2

@@ 3.4 Evaluations with Fine-Tuned LLMs
To enhance the model’s adaptability and responsiveness to our specific task without extensively altering its pre-trained weights, we incorporated the Low-Rank Adaptation (LoRA, [34]) technique. LoRA allows us to dynamically adjust the model’s behavior by introducing trainable parameters that focus on specific parts of the model, such as attention and feed-forward layers. The trainable parameters are around 3.04% for Gemma-2B and 2.29% for Gemma-7B.
@@
Once the model has been thoroughly trained and validated, we apply it to the entire dataset, not limiting its use to only the samples involved in the training process. We then utilize the outcomes from this comprehensive application for subsequent validation. Due to privacy restrictions, we are not permitted to use the Private HNC Dataset for model fine-tuning. As a result, we exclusively use the MIMIC-IV dataset for fine-tuning the model.
@@ 4.1 Introduction of Two EHR Datasets
The MIMIC-IV dataset is an extensive, anon-ymized EHR database encompassing data from over 70,000 patients admitted to Beth Israel Deaconess Medical Center between 2008 and 2019 [9]. This period covers more than a decade of intensive care unit (ICU) admissions, making it a pivotal resource for medical research. With 26 tables of detailed patient data including demographics, lab measurements, procedures, medications, ICD codes, and clinical notes, MIMIC-IV offers comprehensive access to real-world clinical data. This database adheres strictly to the Health Insurance Portability and Accountability Act (HIPAA), ensuring patient privacy while supporting a wide range of medical studies. The breadth and depth of MIMIC-IV facilitate large-scale analyses critical for advancing clinical practice and shaping health policy.
We filter for subjects with at least one clinical note containing metastasis-related keywords, resulting in a dataset comprising notes from 13,923 unique subjects across 28,390 hospital admissions. After examining the ICD codes, 8,510 (approximately 61.1%) of these subjects have ICD-9 codes 197, 198, 199, or ICD-10 codes C78, C79, C80, representing metastasis of tumors. Additionally, 18,552 hospital admissions (approximately 65.3%) are associated with these ICD codes.
@@ 4.2 Evaluation Metrics
“Time-sensitive” sensitivity defined on the Private HNC dataset
Creating a definitive gold standard for evaluating experimental outcomes in this study presents a complex challenge, primarily due to the incomplete reflection of diseases through ICD codes alone. However, given the large volume of clinical notes, expert validation would be both labor-intensive and impractical. To address this and ensure a robust comparative analysis, we employ a methodological approach for our Private HNC Dataset centered around the use of ICD-10 codes C78, C79, and C80, which are indicative of metastatic conditions. The evaluation process is illustrated in Figure 2. Guided by expert input, these codes offer a relatively reliable means of determining a patient’s metastasis status. Therefore, we utilize them as a gold standard to evaluate the effectiveness of various methods. However, it is crucial to acknowledge that the absence of these codes does not unequivocally rule out the presence of metastasis in a patient.
For each patient diagnosed with metastasis according to the aforementioned codes, we review their clinical notes over a 20-day period, encompassing 10 days before and after the diagnosis date. This time frame is chosen to capture the most relevant clinical observations surrounding the diagnosis. Within this period, we systematically search for predefined metastasis-related keywords using regex in the clinical notes as a preprocessing step. We then apply both zero-shot and few-shot learning approaches, with and without the above preprocessing step, to assess the model’s ability to classify metastasis accurately.
The classification accuracy of our model—more precisely, its sensitivity—is assessed based on the relative counts of notes. We do not include a “time-sensitive” specificity analysis, as the absence of an ICD code related to metastasis does not guarantee that no mention of metastasis appears in the clinical notes within a specific time frame. Instead, a specificity analysis is conducted using the MIMIC-IV dataset to validate the robustness of our model. To assess sensitivity, we analyze all clinical notes within the specified time frame. If the number of notes classified by the LLM as indicating metastasis exceeds those indicating its absence, the classification is considered correct. Conversely, if non-metastatic notes predominate, the classification is deemed incorrect. Cases with an equal count are classified as inconclusive due to insufficient information. We denote the total number of correct, incorrect, and inconclusive classifications as $N_{c}$, $N_{i}$, and $N_{l}$, respectively. Additionally, if a review period lacks clinical notes altogether, we interpret this as an absence of relevant information rather than a misclassification by the LLM, and these instances are consequently excluded from our analysis. We use $P_{c}=\frac{N_{c}}{N_{c}+N_{i}+N_{l}}$ to denote the proportion of correct classifications out of all evaluated cases. Similarly, we use $P_{i}=\frac{N_{i}}{N_{c}+N_{i}+N_{l}}$ and $P_{l}=\frac{N_{l}}{N_{c}+N_{i}+N_{l}}$ to denote the proportions of incorrect and inconclusive classifications, respectively.
@@
For the MIMIC-IV dataset, the absence of timestamps for specific ICD codes prevents us from using the same evaluative approach as in the Private HNC Dataset. As a result, we evaluate the results on a per-individual or per-hospital-admission basis. The per-individual evaluation is also depicted in Figure 2. Our process begins by excluding all clinical notes that do not contain keywords related to metastasis. We then apply both regex-based and RAG-based methods to extract metastasis-related sentences for each patient, as preprocessing. Next, we deploy zero-shot and few-shot LLMs, with and without the preprocessing procedure mentioned above, to analyze the remaining notes. We also test various LLMs, such as Gemma-7B-it and the domain-specific LLaMA-2-7B-Chat-Med and Bio-Medical-LLaMA-3-8B, to broaden the evaluation. To ensure robustness, we perform ten additional runs on a subset of the dataset, evaluated exclusively on a per-individual basis. Additionally, we use fine-tuned LLMs to classify the clinical notes, allowing for a comprehensive comparison of different approaches.
@@
We begin by outlining our evaluation method on a per-individual basis. For each patient associated with ICD-9 codes 197, 198, 199, or ICD-10 codes C78, C79, or C80, we conduct a detailed analysis by tallying instances where the LLMs have classified clinical notes as indicative of metastasis. A patient is considered correctly classified if the number of notes flagged as metastatic exceeds those identified as non-metastatic. This method provides a focused evaluation of metastatic conditions within the dataset. To simplify the notation, we let $M_{LLM}^{i}$ and $N_{LLM}^{i}$ denote the group of patients classified as having metastasis and non-metastasis, respectively, by the LLM. Similarly, we let $M_{ICD}^{i}$ and $N_{ICD}^{i}$ represent the group of patients with and without metastasis based on the ICD codes, respectively. We then calculate $P_{c}^{i}=\frac{\#(M_{LLM}^{i}\cap M_{ICD}^{i})}{\#(M_{ICD}^{i})}$ to denote the proportion of patients correctly identified as having metastasis by the LLM out of all patients with a metastatic diagnosis according to ICD codes, where #()#\#($\cdot$)# ( ⋅ ) denotes the cardinality (number of elements) of a set.
@@
We extend this approach to individual hospital admissions, allowing for a more granular assessment of the LLMs’ performance across different patient encounters. We use $M_{LLM}^{ha}$ and $N_{LLM}^{ha}$ to denote the group of hospital admissions classified as having metastasis and non-metastasis, respectively, by the LLM. Similarly, $M_{ICD}^{ha}$ and $N_{ICD}^{ha}$ represent the group of admissions with and without metastasis based on the ICD codes, respectively. We then calculate $P_{c}^{ha}=\frac{\#(M_{LLM}^{ha}\cap M_{ICD}^{ha})}{\#(M_{ICD}^{ha})}$ to denote the proportion of hospital admissions correctly identified as having metastasis by the LLM out of all admissions diagnosed with metastasis according to ICD codes.
@@
The analysis above serves as a measure of “so-called” sensitivity. It is crucial to recognize that even an ideal model is unlikely to achieve 100% sensitivity, as clinicians may not always explicitly document metastasis in clinical notes, despite assigning a corresponding ICD code. Nevertheless, the observed trends offer valuable insights into the model’s performance and its effectiveness in these complex scenarios. Conversely, the absence of an ICD code does not necessarily indicate that metastasis is not mentioned in the notes. Ideally, for patients lacking the relevant ICD codes, the LLMs should more frequently classify them as either not having metastasis or lacking sufficient information. To assess this, we calculate the “so-called” specificity as the proportion of patients without the relevant ICD codes who are correctly classified by the LLMs as non-metastatic or inconclusive, represented by #$\frac{\#(N_{LLM}^{i}\cap N_{ICD}^{i})}{\#(N_{ICD}^{i})}$ for individual patients and #$\frac{\#(N_{LLM}^{ha}\cap N_{ICD}^{ha})}{\#(N_{ICD}^{ha})}$ for hospital admissions. We present the per-individual evaluation flowchart in Figure 2.
@@ 4.3 Results
Figure 3(a) presents a visual summary of the experimental results on our Private HNC Dataset, while the complete numerical results are detailed in Table 1. The results from our investigations substantiate the efficacy of the preprocessing steps integrated into our framework. The results clearly indicate that the regex-based preprocessing significantly enhances the classification accuracy across various modeling paradigms, encompassing both zero-shot and few-shot learning approaches. This enhancement is particularly crucial when employing compact, locally-operated LLMs for processing EHR datasets. Such models, due to their smaller size relative to their more extensive counterparts, often face computational and memory constraints that can limit their effectiveness.
By mitigating these challenges through our targeted preprocessing steps, we are able to considerably improve the utility of smaller LLMs in high-privacy, local computational environments, thereby enhancing their applicability and effectiveness in real-world clinical settings. These findings not only advance our understanding of model scalability and efficiency but also open avenues for further research into optimizing LLM frameworks for specialized data-intensive tasks in constrained environments.
(a) Classification Results for Our Private HNC Dataset.
(b) Classification Results for the MIMIC-IV Dataset.
Figure 3: Comparison of classification results for the Private HNC dataset and MIMIC-IV dataset under different preprocessing conditions and LLM settings. (a) A bar chart displaying classification results for the Private HNC dataset under different preprocessing conditions using Gemma-7B-it. The x-axis represents processing methods (Regex and No Preprocessing), while the y-axis represents the proportion of correctly classified instances. The chart includes three shot settings (Zero-shot, Three-shot, Six-shot) across different time ranges (20 days, 30 days, and 40 days). (b) A bar chart illustrating LLM classification results for the MIMIC-IV dataset under different preprocessing conditions. The x-axis represents processing methods (Regex and No Preprocessing), and the y-axis represents the proportion of correctly classified instances. The chart compares classification accuracy across various LLM configurations, including Zero-shot, Three-shot, and Six-shot learning using Gemma-7B-it, as well as two classification models fine-tuned on Gemma-2B-it and Gemma-7B-it. Each column corresponds to a specific method, while rows distinguish between Patient and Hospital Admission categories.
The results presented in Figure 3(b), with the complete numerical details available in Table 2, underscore the significant impact of regex-based preprocessing on the performance of LLMs in analyzing clinical notes from the MIMIC-IV dataset. This improvement is observed consistently across both zero-shot and few-shot learning settings. Preprocessing leads to significant improvements in classification accuracy across various learning settings. In zero-shot learning on a per-patient basis, preprocessing results in a substantial increase in accuracy compared to models without preprocessing. Similarly, both few-shot settings show large performance gains when preprocessing is applied. This pattern holds true in evaluations based on hospital admissions, where the preprocessed zero-shot model performs far better than the model without preprocessing. The same trend is observed in few-shot models, with preprocessing consistently yielding significantly higher accuracy.
Another notable trend, similar to our findings with the private dataset, is that the model’s performance tends to decline as the number of provided examples increases. This is likely because longer prompts with more examples make it difficult for relatively small models to focus on the key information in the clinical notes. We also observe that with a limited number of samples, the combination of preprocessing and zero-shot learning outperforms fine-tuning the model on those samples before classification. This suggests that preprocessing can be more effective in guiding the model to relevant information than relying on a small amount of fine-tuning data.
Repeated evaluations on subset data yielded results that aligned closely with those obtained from the full dataset. As shown in Figure 4(a), with detailed numerical outcomes provided in Table 4, the chosen preprocessing methods consistently improved model performance across all evaluated LLMs. Note that the previous results focus solely on the “so-called” sensitivity (i.e., power) of the method. Table 3 in the Appendix provides further validation by assessing the “so-called” specificity. This table shows that all methods perform well in terms of specificity. As a result, the primary emphasis should be on sensitivity, where the preprocessed dataset outperforms the others.
(a) Comparison of Different LLMs.
(b) Comparison of Different Preprocessing Approaches.
Figure 4: Comparison of LLMs and preprocessing approaches under different shot settings based on the MIMIC-IV dataset. (a) A bar chart showing the comparison of three different language models (Gemma-7B-it, LLaMA-2-7B-Chat-Med, and Bio-Medical-LLaMA-3-8B) across three shot settings: Zero-shot, Three-shot, and Six-shot. The x-axis represents preprocessing methods, including Regex and No Preprocessing. The y-axis represents Sensitivity. Each row represents a different language model, while each facet represents a different shot setting. Bars are color-coded for different preprocessing methods, and error bars indicate the range of the results. (b) A bar chart comparing different preprocessing approaches under three kinds of shot settings (Zero-shot, Three-shot, Six-shot) for Gemma-7B-it. The x-axis represents preprocessing methods, including Regex, RAG (Top 2), RAG (Top 3), RAG (Top 6), and No Preprocessing. The y-axis represents Sensitivity. Each shot setting is displayed as a separate facet, and error bars indicate the range of the results.
Regex-based preprocessing outperforms RAG on the MIMIC-IV dataset.
Figure 4(b), with detailed numerical results shown in Table 5, presents the sensitivity results comparing the regex-preprocessed, RAG-preprocessed, and no-processing approaches. Across zero-shot, three-shot, and six-shot settings, both regex and RAG exhibit strong overall performance with good consistency (as indicated by the relatively narrow ranges of sensitivity). When using RAG, retrieving the top two sentences yields the highest mean sensitivity; performance declines considerably when retrieving six sentences, likely due to additional irrelevant content that reduces accuracy. Introducing a three-shot scenario further enhances RAG’s comprehension when retrieving the top two relevant sentences, but it still falls short of regex—particularly in the zero-shot setting, where regex achieves its best performance.
Additionally, the RAG approach presents several challenges. One major issue is its reliance on a distance parameter to measure the similarity between the query and candidate sentences for ranking or filtering. Determining an optimal distance threshold is non-trivial, as it lacks a universal reference and varies depending on different keywords and text lengths. Among the tested settings, retrieving the top two sentences yields the best performance, demonstrating RAG’s ability to extract relevant information. However, it still underperforms the regex model. A possible reason is that retrieving only two sentences may omit crucial context, while longer text inputs require additional examples to help the model generate more structured outputs. Conversely, retrieving six sentences introduces excessive irrelevant content, increasing input length and leading to reduced accuracy. Moreover, as the number of retrieved sentences grows, computational overhead rises significantly, further complicating optimization. These challenges make it difficult to establish an optimal threshold for RAG-based preprocessing. In contrast, regex remains a more reliable and broadly applicable method, particularly for smaller LLMs, as it eliminates the need for similarity measures and manual tuning of retrieval parameters.
Summary
All methods demonstrate high specificity and accurately identify non-metastatic cases, the key advantage lies in the models’ sensitivity. The overall conclusion from these results is that preprocessing, which extracts and highlights critical information from clinical notes, significantly improves LLM performance. This paper presents two examples to demonstrate the effectiveness of preprocessing, while other NLP methods may also be applied to the preprocessing step. Regex, guided by predefined references, enhances model performance by extracting precise information from clinical notes. RAG, on the other hand, remains effective in the absence of a universal reference, demonstrating adaptability across various types of unstructured clinical data. This makes preprocessing an essential step when using smaller LLMs on unstructured and complex datasets like EHRs. However, as the number of examples (shots) increases, the performance tends to decline, probably due to input length limitations in smaller models. This highlights the importance of preprocessing for maintaining model efficiency in resource-constrained environments. Furthermore, the findings suggest that when ground truth data is scarce, combining preprocessing with zero-shot learning may be more effective than fine-tuning the model with a limited number of samples. This indicates that in situations where annotation is labor-intensive, particularly for EHR datasets, using preprocessing with zero-shot learning is a more efficient and accurate approach than manual annotation.
5 Discussion
In this study, we introduce a novel compact LLM framework tailored for local deployment on EHR data, addressing significant challenges such as privacy concerns and computational limitations in healthcare settings. Our approach significantly supercharges the performance of smaller LLMs by integrating smart preprocessing techniques, including but not limited to, regex and RAG. This preprocessing not only reduces the data’s complexity before it reaches the LLM but also ensures that the models focus on the most relevant information, thereby improving accuracy and efficiency. Moreover, preprocessing serves as a form of data quality control, helping to refine and standardize input data, ultimately leading to more reliable and interpretable model outputs.
Using “metastasis” as an example, we demonstrate the importance of preprocessing on the Private HNC dataset and MIMIC-IV dataset. The results of our experiments underscore the potential of compact LLMs to perform robustly even under constraints, consistent with previous research that emphasizes the scalability and adaptability of LLMs in various domains. However, unlike previous studies that primarily focus on maximizing model performance without regard for computational efficiency, our work specifically targets the balance between performance and resource utilization, which is critical in many practical healthcare applications. Furthermore, in addition to “metastasis,” we manually assessed the framework’s performance using other keywords, including “Programmed Death-Ligand 1 (PDL1),” “Human Papillomavirus (HPV),” and “Combined Positive Score (CPS).” Although our framework produces reasonably encouraging results, the lack of a definitive gold standard in ICD codes and other resources precludes us from presenting these outcomes in the paper.
Our study provides practical insights into the local deployment of LLMs, particularly for smaller-scale models. Notably, all the compact LLMs we evaluated benefit significantly from preprocessing, highlighting its effectiveness in optimizing performance. Therefore, we recommend prioritizing preprocessing techniques to enhance model accuracy rather than relying solely on fine-tuning. Furthermore, when some domain knowledge is available, simple methods like regex can be highly effective, providing a fast and efficient solution. Notably, this knowledge often does not require much domain expertise—identifying relevant synonyms is usually sufficient, a task that can be easily handled using a free AI tool like ChatGPT. In cases where domain expertise is truly necessary, our study demonstrates that RAG can be a viable alternative. However, while RAG can still yield useful results, it tends to be less effective than regex and requires more computational resources. An additional advantage we did not emphasize earlier is the impact of preprocessing on processing time. Although it introduces an extra step, preprocessing significantly reduces input length, thereby accelerating overall performance. Both regex-based and RAG-based filtering take less than 0.3 seconds per record, with subsequent LLM inference on preprocessed records requiring under 1 second. In contrast, inference on a single unprocessed record averages around 10 seconds. For a dataset exceeding 80,000 clinical notes, this translates to a processing time of 1–2 days with preprocessing, compared to 1–2 weeks without it, marking a substantial improvement in efficiency.
While our findings are promising, our study has certain limitations. The framework’s reliance on regex-based preprocessing may not fully capture the complexities of clinical language, potentially overlooking subtle yet clinically relevant information that falls outside predefined patterns. Although RAG-based methods can help address this limitation, their effectiveness can diminish without optimizing parameters, such as the number of retrieved sentences. As a potential future direction, investigating adaptive thresholding could enhance RAG’s effectiveness. Additionally, the effectiveness of our framework depends on the quality and diversity of the input data, which can vary across different healthcare settings, although we posit that preprocessing will consistently improve results compared to non-preprocessing across all scenarios. Future research could explore the integration of more advanced NLP techniques, such as context-aware embeddings or deep learning-based entity recognition, to further enhance the accuracy and adaptability of the preprocessing step. Moreover, leveraging existing biomedical databases or knowledge graphs could facilitate more advanced preprocessing, enriching the model’s understanding of clinical concepts and relationships. Finally, assessing the framework across a broader range of diseases and clinical conditions would provide deeper insights into its generalizability and robustness.
Appendix A Additional Tables
Table 1 presents the classification sensitivity results for the Private HNC dataset, comparing different preprocessing settings and shot settings. Table 2 summarizes LLM classification sensitivity performance on the MIMIC-IV dataset, evaluating the impact of preprocessing under different shot settings. Table 3 presents the specificity analysis of LLM classification results on the MIMIC-IV dataset. The table compares preprocessed and non-preprocessed data across various shot settings. Table 4 provides a detailed breakdown of LLM performance under different shot settings, specifically comparing Gemma-7B-it, and two LLaMa models (LLaMA-2-7B-Chat-Med, and Bio-Medical-LLaMA-3-8B). The table quantifies differences in classification accuracy across multiple preprocessing conditions. Table 5 compares the effectiveness of different preprocessing approaches, including regex and RAG, under varying shot settings.